OpenStudy (anonymous):

First person 2 give me the right answer gets a badge and Ill fan them Which conjunction or disjunction is equivalent to the open sentence |a – 7| ≤ 5? A. a – 7 ≤ 5 or a – 7 ≥ – 5 B. a – 7 ≥ 5 and a – 7 ≤ – 5 C. a – 7 ≥ 5 or a – 7 ≤ – 5 D. a – 7 ≤ 5 and a – 7 ≥ – 5

OpenStudy (anonymous):

Recall that \(|a|\leq b \implies -b \leq a\leq b\). Hence \(|a-7|\leq 5 \implies -5\leq a-7\leq 5\) How would you write this as a conjunction/disjunction? Can you take things from here? :-)